OpenStudy (anonymous):

List the possible rational zeros of f, then use a graphing utility to determine which of those are actual zeros of f. f(x) = x^3 + 4x^2 -21x

OpenStudy (jdoe0001):

well, I assume you've covered the rational root test

OpenStudy (jdoe0001):

keep in mind that \(\bf f(x) = x^3 + 4x^2 -21x \implies 0= x^3 + 4x^2 -21x\\ \implies 0 = x(x^2 + 4x -21)\)

OpenStudy (anonymous):

how do I find out the possible zeros?